OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6

OpenStudy (lgbasallote):

first perform the distributive prperty. what do you get?

mathslover (mathslover):

first of all what is 3(x+1) ?

OpenStudy (anonymous):

honestly i don't know. i need a tutor :( and i know that you guys giving me the answer is not going to help ''/ but can u break down the steps so i can get it

OpenStudy (lgbasallote):

we werent giving answers heh. that's against the rules of the site :D

mathslover (mathslover):

ok so 3(x+1)=3(x)+3(1)

mathslover (mathslover):

hence we get 3x+3=-2(x-1)+6 right ?

OpenStudy (lgbasallote):

so like we said the first step is distributing 3 into (x+1)